OpenStudy (anonymous):

what is the simplified form of x+1/x^2+x-6 divided by x^2+5x+4/x-2

OpenStudy (anonymous):

x+1/x^2+x-6 divided by x^2+5x+4/x-2 \[\frac{ x + 1 }{ x^2 + x - 6 }\div \frac{ x^2 + 5x + 4 }{ x - 2 } = \frac{ x + 1 }{ x^2 + x - 6 }\times \frac{ x-2 }{ x^2 + 5x + 4 }\]

OpenStudy (anonymous):

\[\[\frac{ x + 1 }{ x^2 + x - 6 }\div \frac{ x^2 + 5x + 4 }{ x - 2 } = \frac{ x + 1 }{ (x -2)(x - 3) }\times \frac{ x-2 }{ (x + 1)(x + 4)}\]\]

OpenStudy (anonymous):

\[\frac{ x + 1 }{ x^2 + x - 6 }\div \frac{ x^2 + 5x + 4 }{ x - 2 } = \frac{1 }{ (x -3)(x + 4) }\]

OpenStudy (anonymous):

\[\frac{ x + 1 }{ x^2 + x - 6 }\div \frac{ x^2 + 5x + 4 }{ x - 2 } = \frac{1 }{ x^2 + x - 12 }\]